OpenStudy (anonymous):

What is the remainder when (3x4 + 2x3 - x2 + 2x - 9) ÷ (x + 2)

OpenStudy (anonymous):

0 5 10 15

OpenStudy (anonymous):

15

OpenStudy (jdoe0001):

\(\bf (3x^4 + 2x^3 - x^2 + 2x - 9) \div (x + 2) \\ \quad \\ \textit{remainder theorem, remainder of }f(x)\ is \\ \quad \\ x+2=0\implies x={\color{red}{ -2}}\qquad remainder\implies f({\color{red}{ -2}})\)

OpenStudy (anonymous):

(3x4+2x3-x^2+2x-9)/(x-2) 12+6-x+2x-9)/(x-2)) 18-3x-9)/(x-) 9-3x/x-2) 9+2-3x/x-2+2 11-3x/x x=-3 11/-3= -3.66 -4 I might be wrong but this is how I worked it out. Check with others. I am new to thing

OpenStudy (anonymous):

This type of problem.